# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v1

@@ 7. Relations to arithmetic geometry
In fact, this conjecture is equivalent to the convergence in Conjecture 1.5, as follows readily from the arithmetic Hilbert–Samuel theorem (by the argument in the proof of [3, Thm 5.3]). One drawback of Conjecture 1.5 – from the arithmetic point of view – is that left hand side in formula 7.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).). This can be remedied if the quantized Futaki character vanishes for $k$ sufficiently large (Section 5.5) and the Lie algebra $\mathfrak{g}$ of the $\mathcal{G}$ is the complexification of a natural $\mathbb{Z}$-module $\mathfrak{g_{\mathbb{Z}}}$ attached to $(\mathcal{X},\mathcal{D})$:
@@
By Prop 5.14, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 5.15 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.6).
@@ 7.1. Examples
The condition 7.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e. $\mathcal{X}=G/P,$ where $G$ is a semisimple Chevalley group over $\mathbb{Z}$ and $P$ a parabolic subgroup of $G$ [66, Section 2], or if $\mathcal{X}$ is the canonical model over $\mathbb{Z}$ of a toric Fano manifold $X$ [81, Section 2]. By Prop 5.20, the assumptions of Conjecture 7.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish. It should be stressed that in these cases, the arithmetic data that go into the definition of the periods $Z_{N}(\mathcal{X},\mathcal{D})$ is, essentially, combinatorial. One is thus led to ask if $Z_{N}(\mathcal{X},\mathcal{D})$ can be computed explicitly in these cases. For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 6.1 and formula 6.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
@@
(and a similar formula holds when $N=4,$ by formula 6.9). Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 6.12 and formula 6.8 gives, when $N\geq 3,$
@@ 8.1. The free energy of a toric log Fano curve
(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [16]). By formula 3.11, this means that
@@
Note that, by formula 8.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
@@ 8.2. The twisted Ding functional of a toric log Fano manifold
(using in the first equality that $\text{(Log)}_{*}\left(|z|^{-2}\frac{i}{2\pi}dz\wedge d\bar{z}\right)=dx$ and [22, Prop 2.9] in the second equality). For example, when $w=0,$ the destabilizing curve for $\mathcal{D}_{\gamma},$ appearing in Cor 4.1 may, if one drops the requirement that it be in $C^{1}(X),$ be expressed explicitly as follows, by taking $v_{0}=0$ in formula 3.35 (which amounts to taking $\phi_{0}(x)\coloneqq|x|)$:
@@
A simple direct calculation gives $\mathcal{D}_{\gamma}(u_{t})=(1/2-\frac{\gamma-1}{\gamma})t+O(1),$ tending to -$\infty$ as $t\rightarrow\infty$ if $\gamma>2.$ This shows that $u_{t}$ is a destabilizing curve for $\mathcal{D}_{\gamma}$ (which provides an alternative direct proof of the destabilizing result in the proof of Cor 4.1). A similar argument applies when $w\in]0,1/2],$ by multiplying the formula for $\phi_{t}(x)$ above by $V/2.$

### v2

@@ 8. Relations to arithmetic geometry
In fact, this conjecture is equivalent to the convergence in Conjecture 1.6, as follows readily from the arithmetic Hilbert–Samuel theorem (by the argument in the proof of [3, Thm 5.3]). One drawback of Conjecture 1.6 – from the arithmetic point of view – is that left hand side in formula 8.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).). This can be remedied if the quantized Futaki character vanishes for $k$ sufficiently large (Section 6.5) and the Lie algebra $\mathfrak{g}$ of the $\mathcal{G}$ is the complexification of a natural $\mathbb{Z}$-module $\mathfrak{g_{\mathbb{Z}}}$ attached to $(\mathcal{X},\mathcal{D})$:
@@
By Prop 6.14, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 6.15 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.7).
@@ 8.1. Examples
The condition 8.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e. $\mathcal{X}=G/P,$ where $G$ is a semisimple Chevalley group over $\mathbb{Z}$ and $P$ a parabolic subgroup of $G$ [66, Section 2], or if $\mathcal{X}$ is the canonical model over $\mathbb{Z}$ of a toric Fano manifold $X$ [81, Section 2]. By Prop 6.20, the assumptions of Conjecture 8.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish. It should be stressed that in these cases, the arithmetic data that go into the definition of the periods $Z_{N}(\mathcal{X},\mathcal{D})$ is, essentially, combinatorial. One is thus led to ask if $Z_{N}(\mathcal{X},\mathcal{D})$ can be computed explicitly in these cases. For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 7.1 and formula 7.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
@@
(and a similar formula holds when $N=4,$ by formula 7.9). Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 7.12 and formula 7.8 gives, when $N\geq 3,$
@@ 9.1. The free energy of a toric log Fano curve
(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [16]). By formula 3.13, this means that
@@
Note that, by formula 9.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
@@ 9.2. The twisted Ding functional of a toric log Fano manifold
(using in the first equality that $\text{(Log)}_{*}\left(|z|^{-2}\frac{i}{2\pi}dz\wedge d\bar{z}\right)=dx$ and [22, Prop 2.9] in the second equality). For example, when $w=0,$ the destabilizing curve for $\mathcal{D}_{\gamma},$ appearing in Cor 4.2 may, if one drops the requirement that it be in $C^{1}(X),$ be expressed explicitly as follows, by taking $v_{0}=0$ in formula 3.37 (which amounts to taking $\phi_{0}(x)\coloneqq|x|)$:
@@
A simple direct calculation gives $\mathcal{D}_{\gamma}(u_{t})=(1/2-\frac{\gamma-1}{\gamma})t+O(1),$ tending to -$\infty$ as $t\rightarrow\infty$ if $\gamma>2.$ This shows that $u_{t}$ is a destabilizing curve for $\mathcal{D}_{\gamma}$ (which provides an alternative direct proof of the destabilizing result in the proof of Cor 4.2). A similar argument applies when $w\in]0,1/2],$ by multiplying the formula for $\phi_{t}(x)$ above by $V/2.$
9.3. Proof of Lemma 4.1 (normalizations)
For any metric $g$ on $S^{2}$ compatible with the complex structure on $\mathbb{P}^{1}$
$dd^{c}u:=\frac{1}{\pi}\frac{i}{2}\partial\bar{\partial}u=\Delta_{g}u\frac{dV_{g}}{4\pi}$
In particular, if $g=g_{S^{2}}$ then $dV_{g}/4\pi$ has unit-mass and thus coincides with $\omega_{\text{FS}}.$ Hence, the defining formula 4.2 for $\Delta_{\text{FS}}$ forces $\Delta_{\text{FS}}=\Delta_{g_{S^{2}}}.$ Next, by formula 3.12,
$E(\mu)=-\frac{1}{2}\int_{X}u_{\mu}\mu,\,\,\text{where}\,dd^{c}(\frac{u_{\mu}}{V})+\omega_{\text{FS}}=\mu,\,\int u_{\mu}\omega_{\text{FS}}=0.$
Let $G(x,y)$ be the symmetric lsc function on $X^{2}$ defined by $dd^{c}G(,\cdot y)=\omega_{\text{FS}}-\delta_{y},$ where $\delta_{y}$ denotes the Dirac measure at $y$ and $\int_{X}G(\cdot,y)\omega_{\text{FS }}=0.$ Then the previous equation yields
$E(\mu)=\frac{V}{2}\int_{X^{2}}G(x,y)\mu\otimes\mu=(1-w)\int_{X^{2}}G(x,y)\mu\otimes\mu.$
Finally, in the holomorphic coordinate $z$ on $\mathbb{C}\subset\mathbb{P}^{1}$ we have $dd^{c}(\log(1+|z|^{2})=\omega_{\text{FS}}$ and $dd^{c}(\log|z|^{2})=\delta_{0}.$ Thus
$G(z,w)=\log\frac{\left|z-w\right|^{2}}{\left(1+\left|z\right|^{2}\right)\left(1+\left|w\right|^{2}\right)}-C,$
$C:=\int_{\mathbb{C}}\log\frac{\left|z\right|^{2}}{1+\left|z\right|^{2}}\omega_{\text{FS}}=\int_{\mathbb{C}}\log\frac{\left|z\right|^{2}}{1+\left|z\right|^{2}}\frac{1}{\pi}\frac{i}{2}\frac{dz\wedge d\bar{z}}{(1+\left|z\right|^{2})^{2}}=\int_{0}^{\infty}\log\frac{r^{2}}{1+r^{2}}\frac{d(r^{2})}{(1+r^{2})^{2}}=-1$
Finally, using stereographic projection gives the explicit formula for $G(\boldsymbol{x},\boldsymbol{y}).$ Indeed, by $SO(3)$-symmetry, the argument of log above equals $(c\left|\boldsymbol{x}-\boldsymbol{y}\right|)^{2}$ for some $c>0$ and using that the diameter of $S^{2}$ equals $2$ forces $c=1/2.$